OpenStudy (anonymous):

Solve 17-8(x+1)=12-3(x+6)

OpenStudy (anonymous):

17-12 = 8x +8 - 3x -18 5 = 5x -10 5x = 15 x = 3

OpenStudy (anonymous):

right? @emily9102

OpenStudy (anonymous):

Im checking it now

OpenStudy (anonymous):

 @smr_kml ?